OpenStudy (nuccioreggie):

A chef draws cookies randomly from a box containing 6 cookies of the same shape and size. There is 1 chocolate cookie, 3 almond cookies, and 2 butter cookies. He draws 1 cookie and then draws another cookie without replacing the first one. Find the probability of picking 1 almond cookie followed by another almond cookie, and show the equation used. Please EXPLAIN THE ANSWER

OpenStudy (nuccioreggie):

@ganeshie8

OpenStudy (nuccioreggie):

@Loser66

OpenStudy (nuccioreggie):

@Mehek14

OpenStudy (nuccioreggie):

@midhun.madhu1987

Mehek (mehek14):

ok first time 3 almond cookies and 6 total

Mehek (mehek14):

so \(\dfrac{3}{6}=\dfrac{1}{2}\)

OpenStudy (nuccioreggie):

Thank you for coming to help me @Mehek14. (:

Mehek (mehek14):

next you don't replace it that means there are only 2 almond cookies left and 5 cookies in total

Mehek (mehek14):

that means \(\dfrac{2}{5}\)

Mehek (mehek14):

so multiply both of them \(\dfrac{1}{2}*\dfrac{2}{5}\)

OpenStudy (nuccioreggie):

Ok can you explain the answer all together please so i will understand better @Mehek14

Mehek (mehek14):

sure

OpenStudy (nuccioreggie):

Ok ty

Mehek (mehek14):

when you choose the almond cookie for the first time, there are 3 almond cookies and 6 cookies in total so the probability of getting an almond cookie is \(\dfrac{3}{6}\) which can be simplified to \(\dfrac{1}{2}\). when you don't put the almond cookie back, there will be only 2 almond cookies in the box and the total will be one less so the total is 5 and the probability is \(\dfrac{2}{5}\) to find your answer, you have to multiply \(\dfrac{1}{2}*\dfrac{2}{5}\)

OpenStudy (nuccioreggie):

Ok then when i multiply that i get my answer?

Mehek (mehek14):

\(\dfrac{2}{10}=\dfrac{1}{5}\)

OpenStudy (nuccioreggie):

@Mehek14. Ok im confused whats the last two fraction for i thought u said i have to muliply those fractions to get me final answer

Mehek (mehek14):

yes 1*2 = 2 2 * 5 = 10 \(\dfrac{2}{10}\)

Mehek (mehek14):

simplify to 1/5

OpenStudy (nuccioreggie):

And 2/10 is the final answer @Mehek14

Mehek (mehek14):

simplified to 1/5

Mehek (mehek14):

so 1/5 is the answer